# Unified approach to the resources of tensor network and stabilizer simulations

### v2

@@ D.2 BKE bounds CBC and BBC
$H_{CBC,\alpha}\left(|O\rangle\right)$ $\geq n-H_{FBKE,1/2}\left(|O\rangle\right),$
$H_{BBC,\alpha}\left(|O\rangle\right)$ $\geq n-H_{BKE,1/2}\left(|O\rangle\right).$ (70)
@@
through which, for both $0\leq\alpha<1$ and $\alpha>1$, we can bound $H_{CBC,\alpha}(|O\rangle)$
$H_{CBC,\alpha}(|O\rangle)$ $=\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2\alpha}\right)$ (75)
@@
$=n-H_{FBKE,1/2}(|O\rangle)+\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2}\right)$
$=n-H_{FBKE,1/2}(|O\rangle),$
where in the second line from the last, we use the definition $H_{FBKE,1/2}(|O\rangle)=H_{1/2}\left(\{\lambda_{i}^{2}\}\right)=2\log\left(\sum_{i}\lambda_{i}\right)$ and in the last line, we use the property $\sum_{ij}|o_{ij}|^{2}=1$. By taking the limit $\alpha\rightarrow 1$, Rényi entropy becomes Shannon entropy, and the same result holds.
@@
By definition (45), we have $c_{i}=2^{-n/2}\text{Tr}(P_{i}O)$. Then for both $0\leq\alpha<1$ and $\alpha>1$,due to the different signal of $(1-\alpha)^{-1}$, $H_{BBC,\alpha}(|O\rangle)$ can always be lower bounded as
$H_{BBC,\alpha}(|O\rangle)$ $=\frac{1}{1-\alpha}\log\left(\sum_{i}|b_{i}|^{2\alpha}\right)=\frac{1}{1-\alpha}\log\left(2^{-n\alpha}\sum_{i}|\mathrm{Tr}(P_{i}O)|^{2\alpha}\right)$ (79)
@@
$=\frac{-n\alpha}{1-\alpha}-H_{BKE,1/2}(|O\rangle)+\frac{n}{1-\alpha}$
$=n-H_{BKE,1/2}(|O\rangle),$
where in the second line from the last, we use the definition $H_{BKE,1/2}(|O\rangle)=H_{1/2}\left(\{\gamma_{i}^{2}\}\right)=2\log\left(\sum_{i}\gamma_{i}\right)$ and the property
@@
In the following box, we summarize the results from Theorem 6-7 and Lemma 2.
Summary of relations around CBC, BBC, and BKE. $H_{CBC,\alpha}(|O\rangle)+H_{BBC,\beta}(|O\rangle)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2,$ $H_{CBC,\alpha}\left(|O\rangle\right)\geq n-H_{FBKE,1/2}\left(|O\rangle\right),$ $H_{BBC,\alpha}\left(|O\rangle\right)\geq n-H_{BKE,1/2}\left(|O\rangle\right),$ $H_{BKE,\alpha}(|O\rangle)+H_{FBKE,\beta}(|O\rangle)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$
@@
$H_{CBC,\alpha}\left(|O\rangle\right)+H_{BBC,\alpha}\left(|O\rangle\right)=2n-\left(H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,1/2}\left(|O\rangle\right)\right).$ (81)
@@
$G_{zero,\alpha}(|O\rangle)\leq\min\left\{n-H_{FBKE,1/2}\left(|O\rangle\right),n-H_{BKE,1/2}\left(|O\rangle\right)\right\}-H_{SE,\alpha}(|O\rangle),\quad O\in\mathcal{T}_{em}.$ (82)
@@ A BKE-matching operator $O$ has minimal CBC-BBC uncertainty
$H_{CBC,\alpha}\left(|O\rangle\right)+H_{BBC,\alpha}\left(|O\rangle\right)=n,\quad O\in\mathcal{T}_{em}.$ (83)
@@
$H_{CBC,\alpha}\left(|O\rangle\right)+H_{BBC,\alpha}\left(|O\rangle\right)=2n-\left(H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,1/2}\left(|O\rangle\right)\right)\geq n,$ (84)
@@
$H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,1/2}\left(|O\rangle\right)\geq H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,\infty}\left(|O\rangle\right)\geq n,$ (85)
@@
$H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,1/2}\left(|O\rangle\right)$ $=n$ (86)
$H_{CBC,\alpha}\left(|O\rangle\right)+H_{BBC,\alpha}\left(|O\rangle\right)$ $=n.$ (87)
@@
Given an BKE-matching operator $O$ with bra-ket entanglement $H_{BKE,1/2}({O})=r\leq n$. If circuit $U$ composed of Clifford and $T$ gates satisfied $H_{SE,\alpha}(U^{\dagger}OU)\geq\eta$, then the $T$ gate number within $U$ satisfied $N_{T}\geq\eta-(n-r)$.
@@
$H_{BKE,1/2}\left(|O\rangle\right)+H_{FBKE,1/2}\left(|O\rangle\right)$ $=n,$ (88)
$H_{CBC,\alpha}\left(|O\rangle\right)+H_{BBC,\alpha}\left(|O\rangle\right)$ $=n,$ (89)
@@
$n-H_{BKE,1/2}\left(|O\rangle\right)$ $\leq H_{BBC,\alpha}\left(|O\rangle\right)$ (90)
$=n-H_{CBC,\alpha}\left(|O\rangle\right)$ (91)
$\leq H_{FBKE,1/2}\left(|O\rangle\right)$ (92)
$=n-H_{BKE,1/2}\left(|O\rangle\right),$ (93)
which means the identity $H_{BBC,\alpha}\left(|O\rangle\right)=n-H_{BKE,1/2}\left(|O\rangle\right)=n-r$. Then noted that
$H_{SE,\alpha}(U^{\dagger}OU)\leq H_{BBC,\alpha}(U^{\dagger}OU)\leq H_{BBC,0}\left(U\otimes U^{*}|O\rangle\right)\leq N_{T}+H_{BBC,0}(|O\rangle).$ (94)
$N_{T}$ $\geq H_{SE,\alpha}(U^{\dagger}OU)-H_{BBC,0}(|O\rangle)$ (95)
@@
Appendix E $\mathcal{F}_{CBC}$ vs $\mathcal{F}_{BBC}$
@@ E.1 Magic-coherence Duality
$\mathcal{F}_{CBC}$ is composed of permutation and diagonal phase gates on the computational basis, and $\mathcal{F}_{BBC}$ are Clifford circuit with also permutation and phase effects on the Pauli basis. However, $\mathcal{F}_{CBC}$ is not merely Clifford circuits with the Pauli basis mapped to the computational basis. The reason is that $U$ can only transform $O$ through $UOU^{\dagger}$, therefore, Clifford circuits acting on a Pauli operator can only introduce phases $\pm 1$. In contrast, $U$ in $\mathcal{F}_{CBC}$ can introduce complex phases for the computational basis. In this sense, $\mathcal{F}_{CBC}$ is larger than $\mathcal{F}_{BBC}$. In fact, we can show that the Clifford structure can be defined in $\mathcal{F}_{CBC}$ and is a subset of $\mathcal{F}_{CBC}$.
@@ E.2 $CCX$ vs $T$
Figure 4: The same setting as Fig 2 except for $U$ in $\mathcal{F}_{CBC}$, we construct them randomly from $\{T,S,CX\}$ instead of $\{CCX,T,S,CX\}$.
@@
In Fig. 4, we can find that when $U$ in $\mathcal{F}_{CBC}$ is made up of $\{T,S,CX\}$, SE cannot approach the upper bound. The reason is that $\{T,S,CX\}$ can only cover a subset of $\mathcal{F}_{CBC}$. For example, $\{T,S,CX\}$ cannot construct $CCX$ [52]. Therefore, we also introduce $CCX$ into the gate set and find that $CCX$ alone without $T$ can already approach the upper bound well as shown in Fig. 2. The reason can be understood from Section E.1, where we show $\{CCX,CX,S\}$ forms the Clifford structure on computational basis, making a pretty good exploration on $\mathcal{F}_{CBC}$ [53].
Appendix F Examples of exponential advantages of operator stabilizer formalism over the standard stabilizer formalism
F.1 Magic-IQP circuit
We consider the simulation of $\mathrm{Tr}(X^{\otimes n}UX^{\otimes n}U^{\dagger})$ with $U=T^{\otimes n}U_{c}T^{\otimes n}$. This is a Pauli version of IQP circuits [54]. In IQP circuit, the amplitude $\langle 0|^{\otimes n}H^{\otimes n}DH^{\otimes n}|0\rangle^{\otimes n}$ is known to be GapP-hard [54, 16]. The basic workflow of IQP circuit is to first create an equally superpositioned state $|+\rangle^{\otimes n}$ on the state computational basis (large CBC) by a layer of $H$, then add phases (with magic $T$) by the diagonal circuit $D$ to each basis, and finally project the state onto $|+\rangle^{\otimes n}$. Similarly, the Magic-IQP circuit $U$ first creates large superpositions of the Pauli basis (large BBC), then adds phases ($\pm 1$) and also permutations to the Pauli basis through Clifford $U_{c}$, and finally traces with an operator with also large Pauli superpositions. Naively, either from the number of $T$ gates or the similarity between magic-IQP and IQP, one might expect that $\mathrm{Tr}(O^{\prime}UOU^{\dagger})$ would be hard to simulate classically. However, due to the Hadamard-like action of $T$ on $X$ and $Y$, and $U$ only has $T$ layers at the beginning and the end of the circuit, in OSF, $\mathrm{Tr}(X^{\otimes n}UX^{\otimes n}U^{\dagger})$ is the overlap between two stabilizer states, making magic-IQP efficiently simulatable.
F.2 X-Y preserving circuit
The action of $T$ to $\{I,X,Y,Z\}$ is identical to the control-Hadamard $CH$ gate in vectorization space with 4 Bell basis. While $CH$ itself is in the third level of the Clifford hierarchy [55], in subspace $\{|I\rangle,|Z\rangle\}$ and $\{|X\rangle,|Y\rangle\}$, $CH$ manifests as $I$ and $H$ respectively. Therefore, for example, if we can restrict the dynamics within the $\{X,Y\}$ subspace, the dynamics in vectorization space become totally Clifford and therefore efficiently simulatable by OSF. Let $O$ in $UOU^{\dagger}$ be $X^{\otimes n}$ and $U$ be made up from the gate set $\{T,SWAP,C_{3}\}$ with $C_{3}$ a special three-qubit Clifford gate, as shown in Ref [45], the requirement is satisfied.

### v3

@@ D.2 BKE bounds CBC and BBC
$H_{\mathrm{CBC},\alpha}(O)$ $\geq n-H_{FBKE,1/2}(O),$
$H_{\mathrm{BBC},\alpha}(O)$ $\geq n-H_{BKE,1/2}(O).$ (72)
@@
through which, for both $0\leq\alpha<1$ and $\alpha>1$, we can bound $H_{\mathrm{CBC},\alpha}(O)$
$H_{\mathrm{CBC},\alpha}(O)$ $=\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2\alpha}\right)$ (77)
@@
$=n-H_{FBKE,1/2}(O)+\frac{1}{1-\alpha}\log\left(\sum_{ij}|o_{ij}|^{2}\right)$
$=n-H_{FBKE,1/2}(O),$
where in the second line from the last, we use the definition $H_{FBKE,1/2}(O)=H_{1/2}\left(\{\lambda_{i}^{2}\}\right)=2\log\left(\sum_{i}\lambda_{i}\right)$ and in the last line, we use the property $\sum_{ij}|o_{ij}|^{2}=1$. By taking the limit $\alpha\rightarrow 1$, Rényi entropy becomes Shannon entropy, and the same result holds.
@@
By definition (47), we have $c_{i}=2^{-n/2}\text{Tr}(P_{i}O)$. Then for both $0\leq\alpha<1$ and $\alpha>1$,due to the different signal of $(1-\alpha)^{-1}$, $H_{\mathrm{BBC},\alpha}(O)$ can always be lower bounded as
$H_{\mathrm{BBC},\alpha}(O)$ $=\frac{1}{1-\alpha}\log\left(\sum_{i}|b_{i}|^{2\alpha}\right)=\frac{1}{1-\alpha}\log\left(2^{-n\alpha}\sum_{i}|\mathrm{Tr}(P_{i}O)|^{2\alpha}\right)$ (81)
@@
$=\frac{-n\alpha}{1-\alpha}-H_{BKE,1/2}(O)+\frac{n}{1-\alpha}$
$=n-H_{BKE,1/2}(O),$
where in the second line from the last, we use the definition $H_{BKE,1/2}(O)=H_{1/2}\left(\{\gamma_{i}^{2}\}\right)=2\log\left(\sum_{i}\gamma_{i}\right)$ and the property
@@
In the following box, we summarize the results from Theorem 5-6 and Lemma 3.
Summary of relations around CBC, BBC, and BKE. $H_{\mathrm{CBC},\alpha}(O)+H_{BBC,\beta}(O)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2,$ $H_{\mathrm{CBC},\alpha}(O)\geq n-H_{FBKE,1/2}(O),$ $H_{\mathrm{BBC},\alpha}(O)\geq n-H_{BKE,1/2}(O),$ $H_{BKE,\alpha}(O)+H_{FBKE,\beta}(O)\geq n,\quad\forall~\alpha,\beta\geq\frac{1}{2}\quad\text{and}\quad\frac{1}{\alpha}+\frac{1}{\beta}=2.$
@@
$H_{\mathrm{CBC},\alpha}(O)+H_{\mathrm{BBC},\alpha}(O)=2n-\left(H_{BKE,1/2}(O)+H_{FBKE,1/2}(O)\right).$ (83)
@@
$G_{zero,\alpha}(O)\leq\min\left\{n-H_{FBKE,1/2}(O),n-H_{BKE,1/2}(O)\right\}-H_{SE,\alpha}(O),\quad O\in\mathcal{T}_{em}.$ (84)
@@ A BKE-matching operator $O$ has minimal CBC-BBC uncertainty
$H_{\mathrm{CBC},\alpha}(O)+H_{\mathrm{BBC},\alpha}(O)=n,\quad O\in\mathcal{T}_{em}.$ (85)
@@
$H_{\mathrm{CBC},\alpha}(O)+H_{\mathrm{BBC},\alpha}(O)=2n-\left(H_{BKE,1/2}(O)+H_{FBKE,1/2}(O)\right)\geq n,$ (86)
@@
$H_{BKE,1/2}(O)+H_{FBKE,1/2}(O)\geq H_{BKE,1/2}(O)+H_{FBKE,\infty}(O)\geq n,$ (87)
@@
$H_{BKE,1/2}(O)+H_{FBKE,1/2}(O)$ $=n$ (88)
$H_{\mathrm{CBC},\alpha}(O)+H_{\mathrm{BBC},\alpha}(O)$ $=n.$ (89)
@@
Given an BKE-matching operator $O$ with bra-ket entanglement $H_{BKE,1/2}({O})=r\leq n$. If circuit $U$ composed of $\{H,S,CNOT,T\}$ gates satisfied $H_{SE,\alpha}(U^{\dagger}OU)\geq\eta$, then the $T$ gate number within $U$ satisfied $N_{T}\geq\eta-(n-r)$ and $N_{H}\geq\eta-r$.
@@
$H_{BKE,1/2}(O)+H_{FBKE,1/2}(O)$ $=n,$ (90)
$H_{\mathrm{CBC},\alpha}(O)+H_{\mathrm{BBC},\alpha}(O)$ $=n,$ (91)
@@
$n-H_{BKE,1/2}(O)$ $\leq H_{\mathrm{BBC},\alpha}(O)$ (92)
$=n-H_{\mathrm{CBC},\alpha}(O)$ (93)
$\leq H_{FBKE,1/2}(O)$ (94)
$=n-H_{BKE,1/2}(O),$ (95)
which means the identity $H_{\mathrm{BBC},\alpha}(O)=n-H_{BKE,1/2}(O)=n-r$. Then noted that
$H_{SE,\alpha}(U^{\dagger}OU)\leq H_{\mathrm{BBC},\alpha}(U^{\dagger}OU)\leq H_{BBC,0}(U^{\dagger}OU)\leq N_{T}+H_{BBC,0}(O).$ (96)
$N_{T}$ $\geq H_{SE,\alpha}(U^{\dagger}OU)-H_{BBC,0}(O)$ (97)
@@
Similarly, for the $H$ gates, we conclude that:
$H_{SE,\alpha}(U^{\dagger}OU)\leq H_{\mathrm{CBC},\alpha}(U^{\dagger}OU)\leq H_{BBC,0}(U^{\dagger}OU)\leq 2N_{H}+H_{CBC,0}(O),$ (99)
which leads to the result $N_{H}\geq(1/2)\cdot[H_{SE,\alpha}(U^{\dagger}OU)-r]$ based on the fact that $H_{\mathrm{BBC},\alpha}(O)=r$ for TE-matching operators. ∎
Appendix E $\mathcal{F}_{\mathrm{CBC}}$ vs $\mathcal{F}_{\mathrm{BBC}}$
@@ E.1 Magic-coherence Duality
$\mathcal{F}_{\mathrm{CBC}}$ is composed of permutation and diagonal phase gates on the computational basis, and $\mathcal{F}_{\mathrm{BBC}}$ are Clifford circuit with also permutation and phase effects on the Pauli basis. However, $\mathcal{F}_{\mathrm{CBC}}$ is not merely Clifford circuits with the Pauli basis mapped to the computational basis. The reason is that $U$ can only transform $O$ through $UOU^{\dagger}$, therefore, Clifford circuits acting on a Pauli operator can only introduce phases $\pm 1$. In contrast, $U$ in $\mathcal{F}_{\mathrm{CBC}}$ can introduce complex phases for the computational basis. In this sense, $\mathcal{F}_{\mathrm{CBC}}$ is larger than $\mathcal{F}_{\mathrm{BBC}}$. In fact, we can show that the Clifford structure can be defined in $\mathcal{F}_{\mathrm{CBC}}$ and is a subset of $\mathcal{F}_{\mathrm{CBC}}$.
@@ E.2 $CCX$ vs $T$
Figure 4: The same setting as Fig 3 except for $U$ in $\mathcal{F}_{\mathrm{CBC}}$, we construct them randomly from $\{T,S,CX\}$ instead of $\{CCX,T,S,CX\}$.
@@
In Fig. 4, we numerically find that when $U$ in $\mathcal{F}_{\mathrm{CBC}}$ is made up of $\{T,S,CX\}$, SE cannot approach the upper bound. The simulations were performed using mindquantum package [46, 13]. The reason is that $\{T,S,CX\}$ can only cover a subset of $\mathcal{F}_{\mathrm{CBC}}$. For example, $\{T,S,CX\}$ cannot construct $CCX$ [32]. Therefore, we also introduce $CCX$ into the gate set and find that $CCX$ alone without $T$ can already approach the upper bound well as shown in Fig. 3. The reason can be understood from Section E.1, where we show $\{CCX,CX,S\}$ forms the Clifford structure on computational basis, making a pretty good exploration on $\mathcal{F}_{\mathrm{CBC}}$ [45].
Appendix F Application of BKE-matching operators
We identify three physically motivated settings where BKE-matching operators with intermediate BKE arise naturally, demonstrating that the regime beyond zero BKE is essential for modern quantum simulation tasks.
Marginal probability distributions.
Weak simulation of quantum circuits can be reduced to computing marginal probability distributions [9]. For an $n=(n_{1}+n_{2})$-qubit system, the marginal over $n_{2}$ qubits reads
$p(\mathbf{s})=\mathrm{tr}\!\left[U|0\rangle\!\langle 0|^{\otimes n}U^{\dagger}\cdot\left(\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|\right)\right].$ (118)
The observable $O=\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|$ is a tensor product of single-qubit identity and computational-basis operators, hence a BKE-matching operator with $H_{\mathrm{BKE},1/2}(O)=n_{1}$ and $H_{\mathrm{FBKE},1/2}(O)=n_{2}$, giving $H_{\mathrm{CBC},\alpha}(O)=n_{2}$ and $H_{\mathrm{BBC},\alpha}(O)=n_{1}$. Applying Lemma 8, if a Clifford+$T$ circuit $U$ achieves $H_{\mathrm{SE},\alpha}(U^{\dagger}OU)\geq\eta$, then the gate counts satisfy
$N_{T}\geq\eta-n_{2},\qquad N_{H}\geq\eta-n_{1}.$ (119)
For the extremal cases, $n_{1}=0$ (amplitude estimation) recovers the purely coherence-limited regime with cost scaling as $2^{N_{H}}$, while $n_{1}=n$ (full distribution) recovers the magic-limited regime with cost $2^{N_{T}}$. At intermediate values $n_{1}\approx n_{2}\approx n/2$, both bounds impose comparable constraints, confirming that neither computational-basis nor Pauli-basis methods alone are efficient.
Mixed-state inputs.
Experimental quantum devices inevitably produce mixed initial states due to imperfect preparation. Consider $n_{1}$ depolarized qubits subject to a Pauli error $P$ and $n_{2}$ well-initialized qubits:
$\rho_{0}=\frac{\mathbb{I}_{n_{1}}+\varepsilon P}{2^{n_{1}}}\otimes|0\rangle\!\langle 0|^{\otimes n_{2}}=2^{-n_{1}}O_{1}+2^{-n_{1}}\varepsilon\,O_{2},$ (120)
where $O_{1}=\mathbb{I}_{n_{1}}\otimes|0\rangle\!\langle 0|^{\otimes n_{2}}$ and $O_{2}=P\otimes|0\rangle\!\langle 0|^{\otimes n_{2}}$. Both $O_{1}$ and $O_{2}$ are BKE-matching operators, being tensor products of single-qubit Pauli and computational-basis operators. This decomposition is exact and requires only two terms, whereas a computational-basis decomposition of $\mathbb{I}_{n_{1}}$ would require $2^{n_{1}}$ terms and a full Pauli decomposition of $|0\rangle\!\langle 0|^{\otimes n_{2}}$ would require $2^{n_{2}}$ terms. The output probability $p(\mathbf{x})=\mathrm{tr}(U\rho_{0}U^{\dagger}\cdot|\mathbf{x}\rangle\!\langle\mathbf{x}|)$ thus reduces to two BKE-matching calculations, each analyzable via Lemma 8 with the BKE determining the coherence-versus-magic resource balance.
Quantum error correction protocols.
Syndrome extraction in fault-tolerant quantum computing combines both of the above settings. Data qubits ($n_{1}$ qubits, subject to logical Pauli errors) and ancilla qubits ($n_{2}$ qubits, initialized to $|0\rangle\!\langle 0|^{\otimes n_{2}}$) undergo a syndrome extraction circuit $U$. The syndrome probability is
$p(\mathbf{x},\mathbf{s})=\mathrm{tr}\!\left[U\rho_{0}U^{\dagger}\cdot\left(\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|\right)\right],$ (121)
where $\rho_{0}=2^{-n_{1}}O_{1}+2^{-n_{1}}\varepsilon\,O_{2}$ as before. This calculation involves the contraction of BKE-matching operators from both the input ($O_{1}$, $O_{2}$) and the output measurement ($\mathbb{I}_{n_{1}}\otimes|\mathbf{s}\rangle\!\langle\mathbf{s}|$), all residing in the intermediate BKE regime. The simulation cost for each term is quantified by Lemma 8, with the BKE values of input and output operators jointly determining whether coherence or magic resources dominate.